# SentenceVAE: Faster, Longer and More Accurate Inference with Next-sentence Prediction for Large Language Models

### v1

@@ 1 Hello, 3 Hello, Hello, Hello,
Experiment
To validate our idea, we first trained individual SentienceVAEs using self supervised methods, demonstrating that multiple tokens in a sentence can be compressed into a single sentence vector by an encoder and restored by a decoder. Subsequently, we grafted the encoder and decoder at the beginning and end of the open-source LLMs, proving that with simple modifications, LLMs can be upgraded to SLLMs and work in sentence embedding space, while preserving PPL and improving inference speed. Meanwhile, by observing the loss curve, we found that SLLMs still follows the Scaling Law.
Experimental Setting
The experiments described in this paper were conducted using either a single 4-card RTX 4090 (24G) or a 4-card A100 (40G) (for SLLM-1.3B) setup, with training carried out in a data parallel distribution.
Base Models: Due to memory and time limitations, we only use the 125m, 350m, and 1.3B models of the OPT series (Zhang et al. 2022) as the base LLMs, and demonstrate the feasibility of extending to larger scale models through Scaling Law (Kaplan et al. 2020).
Dataset: We use the English subset (EN/WebText) of the Wanjuan-1.0 dataset (He et al. 2023) as the training dataset. The number of training samples depends on the number of iterations, with SentenceVAE training samples consisting of approximately 153.6M sentences (approximately 1.7B tokens); The training sample size of SLLM is about 6.4M paragraphs (approximately 5.6B tokens). The validation set consists of 1,000 random sentences or paragraphs that are mutually exclusive with the training set.
Hyperparameters: When training SentenceVAE, we assume each sentence has a maximum of 64 tokens, with a batch size of 128/card, a base learning rate of 1e-7/batch/card, and a maximum of 300K iterations.
For training SLLM, we assume each sentence has a maximum of 64 tokens, each paragraph has a maximum of 64 sentences (due to memory limitations), a batch size of 1/card, a base learning rate of 1e-6/batch/card, and a maximum of 1.6M iterations.
All experiments were conducted using the AdamW optimizer (Loshchilov and Hutter 2017) with AMP enabled (Micikevicius et al. 2017), a weight decay coefficient of 0.01, and a maximum L2 norm of 1 for gradient clipping.
The first 5K iterations utilized a linear learning rate schedule, starting at 1/1000 of the base learning rate. After 5K iterations, a cosine annealing learning rate schedule was employed with a period of 20K iterations. Additionally, all experiments implemented the Exponential Moving Average (EMA) strategy.
Metrics: We use perplexity (PPL) as an evaluation metric. The formula for calculating perplexity for output logits $\boldsymbol{P}$ of Sentence Decoder and ground truth $\boldsymbol{D}_{\text{groud\_truth}}$ is as Eq. 10.
(a) Test loss curve of SVAEs.
(b) Test loss curve of SLLMs.
Figure 3: Scaling Law of SVAEs and SLLMs.
A series of word-level tokens can be represented by a sentence-level embedding with strong robustness
In this experiment, we trained individual SentienceVAEs using a self-supervised approach. If the Sentence Encoder can compress multiple tokens from the input sentence into a single sentence embedding vector and the Sentence Decoder can use this vector to reconstruct the original token sequence, it would validate the feasibility of our theory.
We trained SentenceVAEs with hidden sizes of 768, 1024, and 2048, corresponding to OPT-125M, OPT-350M, and OPT-1.3B, respectively. For each hidden size, we varied the number of block layers to 1, 2, and 4. We evaluated the models using cross-entropy loss and perplexity (PPL) metrics on the validation set. The experimental results are presented in Table 1.
The evaluation results demonstrate that SentenceVAEs fulfill their intended purpose, confirming the theoretical validity. Table 2 presents several test cases. Among these, Samples 1, 2, 3, 5, 6 are convention samples. To assess the model’s robustness, we also included several out-of-distribution samples, highlighted in blue in the table:
@@
  • Sample 4, 7: Data preprocessing involves segmenting sentences based on punctuation marks, such as "," and ".". Consequently, these samples will be split into multiple inputs during both training and practical application. Directly inputting these samples without preprocessing is considered invalid input for the model.
@@
  • Sample 8: This is the pinyin of a famous Chinese saying, not English, so it is illegal input.
From the output results of the model, we can analyze that:
@@
  • In the convention samples, except for Sample 6, the model accurately restored the sentences. The reason for the error in Sample 6 may be related to the bottleneck of the sinusoidal position encoding mechanism (Vaswani et al. 2017).
@@
  • Although Sample 4 contains the illegal input, the model has accurately restored it with only punctuation marks. Sample 7 still maintains the meaning of the sentence. In Sample 8, as the hidden size of the model increases, the model also accurately restores the illegal input “yao yao ling xian”. These phenomena indicate that the model is robust to illegal inputs.
@@
LLMs can work in setence-level embedding space with faster inference speed, more accurate PPL, and longer context, resulting SLLMs
After verifying the feasibility of Sentence level embedding, we grafted the encoder and decoder of SentenceVAE onto the beginning and end of the OPT series models, corresponding to the relationship shown in Table 4.
Model Base Model Grafted SentenceVAE
SLLM-125M-H1 OPT-125M SVAE-768-H1
SLLM-125M-H2 SVAE-768-H2
SLLM-125M-H4 SVAE-768-H4
SLLM-350M-H1 OPT-350M SVAE-1024-H1
SLLM-350M-H2 SVAE-1024-H2
SLLM-350M-H4 SVAE-1024-H4
SLLM-1.3B-H1 OPT-1.3B SVAE-2048-H1
SLLM-1.3B-H2 SVAE-2048-H2
We did not test SLLM-1.3B-H4 due to memory overflow issues encountered during data-parallel distributed training. The benchmark test results of OPTs and SLLMs are shown in Table 3.
Faster inference speed: We assessed the throughput rate of SLLMs using the PyTorch framework (Imambi, Prakash, and Kanagachidambaresan 2021), without incorporating optimization methods such as PagedAttention (Kwon et al. 2023) or acceleration engines like TensorRT-LLM or LMDeploy (Contributors 2023). The throughput rate of the original OPT model served as the baseline for comparison. The experimental results show that SLLMs is 2$\sim$3$\times$ faster on average than LLMs. The larger the number of model parameters, the smaller the proportion of inferencing cost occupied by SentenceVAE, and the more obvious the speed benefit.
More accurate PPL: We observed that the PPL metric for SLLMs was better than that of the baseline OPT models. This improvement is likely due to the SLLM paradigm enabling the model to process natural language at a higher sentence level, leading to enhanced performance. Additionally, the attention computation in the sentence embedding space allows SLLMs to manage shorter relative contexts for texts of the same length, which further contributes to the improved model performance.
Longer context length: The context length that a model can handle is primarily constrained by GPU memory. Under the SLLM framework, multiple original tokens can be compressed into a single token, allowing the model to use less memory when processing contexts of the same length. This effectively increases the context length that the model can handle given the same hardware resources.
@@
By analyzing the loss curve of SLLM during the training process, we observed that it adheres to the Scaling Law (Kaplan et al. 2020), as shown in Fig. 3. This suggests that the SLLM framework can be effectively extended to larger-scale language models through scaling.
@@
In this section, we summarize our work and suggest several potential directions for future development.
Scaling up SLLMs with better structures
In this work, to rapidly validate our ideas, we utilized the classic Transformer architecture (Vaswani et al. 2017) implemented in the PyTorch framework (Imambi, Prakash, and Kanagachidambaresan 2021). Due to constraints in computing power and time, we limited our validation to models ranging from 125M to 1.3B parameters. However, by verifying the Scaling Law (Kaplan et al. 2020), we inferred the applicability of our approach to larger scale models. Numerous state-of-the-art architectural optimizations can be integrated into the SLLM framework, such as Rotational Position Encoding (RoPE) (Su et al. 2024), which may address the order problem observed in Sample 6. Currently, our models are trained solely on English corpora, but there is potential to add multilingual support in the future. In summary, while the SLLM framework is still in its early stages, it holds significant potential for further optimization.
SLLMs assist in hybrid edge-cloud inference
The SLLM framework is inherently a hybrid architecture combining small and large models. By identifying the optimal balance point, the small model (SentenceVAE) and the large model (LLM) can be effectively distributed between the edge and cloud, balancing computational load and enhancing user experience.
SLLMs assist in embodied intelligence
Currently, embodied intelligence based on the LLM Agent paradigm struggles to interact directly with underlying hardware, relying instead on intermediate components to translate high-level LLM instructions into traditional real-time control logic. The primary limitation is the insufficient rate of “token” generation on the edge side. SLLM, however, can process and generate more “tokens” within the same computational and temporal constraints, offering the potential for embodied intelligent large models to interface directly with underlying hardware.
SLLMs assist in multimodal large model
The “frames” in videos and “trunks” in audio are, to some extent, equivalent to “tokens” in language models. Therefore, the SLLM paradigm can similarly be applied to multimodal large models to increase the processing “frame rate”, enhance user experience, and achieve performance comparable to or even surpassing that of GPT-4o.

### v3

@@ 1 Hello, 3 Hello, Hello, Hello,
Figure 3: Scaling Law of (a) SLLMs and (b) SVAEs.
The evaluation results affirm that SentenceVAEs achieve their intended functionality, substantiating their theoretical underpinnings. Table 2 showcases a variety of test cases, including Samples 1, 2, 3, 5, 6, which serve as conventional examples. To evaluate model robustness, several out-of-distribution samples, marked in blue, are also incorporated:
@@
  • Sample 4, 7: data preprocessing entails segmenting sentences based on punctuation marks, such as "," and ".". Consequently, these samples are split into multiple inputs during both training and deployment. Inputting these samples without prior preprocessing constitutes invalid input for the model.
@@
  • Sample 8: this sample is the pinyin transcription of a well-known Chinese saying and is not in English, hence it is considered invalid input.
Analysis of the model’s output reveals the following:
@@
  • In the conventional samples, the model accurately reconstructed the sentences, with the exception of Sample 6. The error in Sample 6 may be attributed to limitations inherent in the sinusoidal position encoding mechanism (Vaswani et al. 2017).
@@
  • Although Sample 4 contains the illegal input, the model has accurately restored it with only error of punctuation marks. Sample 7 still maintains the meaning of the sentence. In Sample 8, as the hidden size of the model increases, the model also accurately restores the illegal input “yao yao ling xian”. These phenomena indicate that the model is robust to illegal inputs.
SLLMs can work in setence-level embedding space with faster inference speed, more accurate PPL, and longer context
Upon validating the practicality of sentence-level embeddings, we integrated the encoder and decoder components from SentenceVAE with the OPT series models. This integration is detailed in Table 3.
Model Base Model Grafted SentenceVAE
SLLM-125M-H1 OPT-125M SVAE-768-H1
SLLM-125M-H2 SVAE-768-H2
SLLM-125M-H4 SVAE-768-H4
SLLM-350M-H1 OPT-350M SVAE-1024-H1
SLLM-350M-H2 SVAE-1024-H2
SLLM-350M-H4 SVAE-1024-H4
SLLM-1.3B-H1 OPT-1.3B SVAE-2048-H1
SLLM-1.3B-H2 SVAE-2048-H2
@@
Due to memory constraints during data-parallel distributed training, we were unable to evaluate the SLLM-1.3B-H4 model. Comparative performance metrics for the OPT and SLLM models are presented in Table 4.
Faster Inference: the throughput of SLLMs, measured using the PyTorch framework (Imambi, Prakash, and Kanagachidambaresan 2021), was assessed without the inclusion of optimization techniques such as PagedAttention (Kwon et al. 2023) or acceleration frameworks like TensorRT-LLM or LMDeploy (Contributors 2023). The baseline for this comparison was the original OPT model’s throughput. Experimental findings indicate that SLLMs achieve an average inference speed that is 2 to 3 times faster than traditional LLMs. Notably, as the model size increases, the relative overhead of SentenceVAE decreases, thereby magnifying the speed advantage.
Higher Accuracy: the PPL scores for SLLMs were superior to those of the baseline OPT models. This enhancement is attributed to the SLLM framework’s ability to process language at a more granular sentence level, thereby boosting overall performance. Moreover, the computation of attention within the sentence embedding space enables SLLMs to handle shorter relative contexts for equivalent text lengths, further enhancing model efficacy.
Longer Context: the maximum context length manageable by a model is generally limited by GPU memory availability. The SLLM framework compresses multiple original tokens into a single token, reducing memory usage for equivalent-length contexts. This compression allows for an extended context capacity within the same hardware constraints.
@@
During the training of SLLMs, an analysis of the loss curve indicates compliance with the Scaling Law, as depicted in Figure 3. This observation supports the hypothesis that the SLLM framework is amenable to effective scaling up to larger language models.
@@
This section encapsulates the contributions of our research and delineates several promising avenues for future exploration.
Scaling up SLLMs with Enhanced Architectures
In our study, we employed the well-established Transformer architecture (Vaswani et al. 2017), implemented using the PyTorch framework (Imambi, Prakash, and Kanagachidambaresan 2021), to expedite the validation of our hypotheses. Constraints related to computational resources and time necessitated that our evaluations were confined to models with parameter sizes ranging from 125M to 1.3B. Nevertheless, by corroborating the Scaling Law, we extrapolated the feasibility of our methodologies to larger-scale models. Future iterations of the SLLM framework could integrate advanced architectural enhancements, such as Rotational Position Encoding (RoPE) (Su et al. 2024), to rectify sequence ordering issues identified in Sample 6. Although our current models are trained exclusively on English-language corpora, extending support to multiple languages could be a beneficial direction. Despite its nascent stage, the SLLM framework exhibits substantial potential for enhancements and broader applicability.
SLLMs in Hybrid Edge-Cloud Inference
The SLLM framework embodies a hybrid architecture that amalgamates small and large models. By pinpointing an optimal balance, the smaller model (SentenceVAE) and the larger model (LLM) can be efficiently allocated across edge and cloud environments, optimizing computational loads and improving user interactions.
SLLMs and Embodied Intelligence
Currently, embodied intelligence that leverages the LLM Agent paradigm faces challenges in direct hardware interactions, relying instead on intermediary mechanisms to translate high-level LLM directives into conventional real-time control logic. A critical bottleneck is the generation rate of “tokens” at the edge. In contrast, the SLLM framework’s enhanced ability to process and produce more “tokens” under identical computational and temporal constraints presents opportunities for direct interfacing of large embodied intelligence models with underlying hardware systems.
SLLMs in Multimodal Large Models
In the context of multimodal large models, the “frames” in video and “trunks” in audio data can be analogized to “tokens” in text-based models. The SLLM method could thus be adapted to augment the processing rates of “frames” in these modalities, potentially elevating user experiences and achieving performance that matches or surpasses current standards.
Acknowledgments
This research was supported by the China National Key R&D Program (2022YFC2808003) and the Natural Science Basic Research Program of Shaanxi (2024JC-YBMS-468).